### v1

# SiDiaC: Sinhala Diachronic Corpus

Abstract
SiDiaC, the first comprehensive Sinhala Diachronic Corpus, covers a historical span from the 5th to the 20th century CE. SiDiaC comprises 58k words across 46 literary works, annotated carefully based on the written date, after filtering based on availability, authorship, copyright compliance, and data attribution. Texts from the National Library of Sri Lanka were digitised using Google Document AI OCR, followed by post-processing to correct formatting and modernise the orthography. The construction of SiDiaC was informed by practices from other corpora, such as FarPaHC, particularly in syntactic annotation and text normalisation strategies, due to the shared characteristics of low-resourced language status. This corpus is categorised based on genres into two layers: primary and secondary. Primary categorisation is binary, classifying each book into Non-Fiction or Fiction, while the secondary categorisation is more specific, grouping texts under Religious, History, Poetry, Language, and Medical genres. Despite challenges including limited access to rare texts and reliance on secondary date sources, SiDiaC serves as a foundational resource for Sinhala NLP, significantly extending the resources available for Sinhala, enabling diachronic studies in lexical change, neologism tracking, historical syntax, and corpus-based lexicography.
SiDiaC: Sinhala Diachronic Corpus
Nevidu Jayatilleke  and Nisansa de Silva Department of Computer Science & Engineering, University of Moratuwa, Sri Lanka {nevidu.25, NisansaDdS}@cse.mrt.ac.lk
1 Introduction
Sola lingua bona est lingua mortua [^1] Latin: The only good language is a dead language.; given that all languages that are in use, evolve with a gradual process of linguistic change over time, proposed to have originated from an initially gestural communication system Corballis (2017). Factors affecting this complex evolution include cultural influences, which drive irregular word meaning shifts through phenomena such as new technologies (e.g., cell to mean cell phone in addition to prison cell) or community-specific vernaculars (e.g., gay to mean homosexual in addition to carefree). These cultural shifts often impact nouns more significantly. Conversely, regular linguistic processes, such as subjectification (e.g., actually shifting from objective to subjective usage) or grammaticalisation (e.g., promise undergoing rich changes), cause more predictable semantic changes and tend to affect verbs, adjectives, and adverbs more readily Hamilton et al. (2016).
The Sinhala language is an Indo-European language, which possesses a rich and diverse literary heritage that has developed over the course of several millennia, with its origins tracing back to between the 3rd and 2nd centuries BCE  de Silva (2025). This language has undergone significant evolution and transformation throughout its history, resulting in the form of modern Sinhala that we engage with today. Sinhala is spoken as L1 by approximately 16 million people, primarily located on the island of Sri Lanka de Silva (2025). The Sinhala script, which is unique to the language, descends from the Indian Brahmi script Fernando (1949); De Mel et al. (2025). Sinhala is classified as a lower-resourced language (Category 02) according to the criteria presented by Ranathunga and de Silva (2022).
In this study, we introduce a novel diachronic Sinhala dataset, SiDiaC [^2] https://github.com/NeviduJ/SiDiaC, which covers the period from 426 CE to 1944 CE. This dataset is based on distinct identifications of written years or specific time frames of the recognised Sinhala literature.
2 Existing Work
The development of historical corpora has garnered significant attention due to its importance beyond the creation of general-purpose corpora. It enables researchers to investigate the evolution of language, taking into account changes in semantics, lexicon, morphology, and syntax. As a result, studies have been conducted to develop diachronic corpora for different languages.
2.1 LatinlSE
McGillivray and Kilgarriff (2013) introduced LatinlSE, a 13-million-word historical Latin corpus developed for the Sketch Engine [^3] https://www.sketchengine.eu/, a leading corpus query tool. Covering an extensive 22-century period from the 2nd Century BCE to the 21st Century CE, LatinlSE is equipped with detailed metadata, including author, title, genre, era, date, and century.
The methodology for creating this corpus involved gathering texts from various online digital libraries, such as LacusCurtius [^4] https://penelope.uchicago.edu/Thayer/E/Roman/Texts/, IntraText [^5] https://www.intratext.com/, and Musisque Deoque [^6] https://www.mqdq.it/. This process ensured a broad classification of genres as prose and poetry, and the texts were converted into a verticalized format while preserving their metadata. A significant aspect of the creation process was the automatic linguistic annotation using advanced NLP tools. This included lemmatisation with the PROIEL [^7] https://www.hf.uio.no/ifikk/english/research/projects/proiel/ project’s morphological analyser, complemented by Quick Latin [^8] https://www.quicklatin.com/ for unrecognised forms. Part-of-Speech (POS) tagging was achieved by training TreeTagger Schmid (1999) on existing Latin treebanks, including the Index Thomisticus Treebank [^9] https://itreebank.marginalia.it/, the Latin Dependency Treebank Bamman and Crane (2006), and the PROIEL project’s Latin treebank Haug and Jøhndal (2008). This training helps disambiguate analyses and assign the most likely lemma and POS to each token in context. This comprehensive dataset allows users to perform sophisticated searches based on lemmas, POS, and context, facilitating the study of shifts in word meanings over time.
2.2 IcePaHC and FarPaHC
The Icelandic Parsed Historical Corpus (IcePaHCRögnvaldsson et al. (2012) is a one-million-word parsed historical corpus of Icelandic, spanning from the late 12th century to the early 21st century. But more relevant to our work in this study is The Faroese Parsed Historical Corpus (FarPaHC), a syntactically annotated corpus of Faroese historical texts, that is presented as a spin-off of IcePaHC. The reason for this relevance is that, according to Ranathunga and de Silva (2022), Faroese also belongs to Category 02, similar to Sinhala. The FarPaHC corpus has 53,000 words.
It’s given that FarPaHC is an extension of IcePaHC; the primary sources included narrative and religious texts that have parallel texts in IcePaHC. A key step in the process was the conversion of all texts to modern spelling using the IceNLP package [^10] https://sourceforge.net/projects/icenlp/ (which includes a tokeniser, POS tagger, and lemmatiser), which was necessary for preprocessing and for facilitating searches. The annotation process involved manually dividing clauses, semi-automatically preprocessing texts with IceNLP and CorpusSearch [^11] https://corpussearch.sourceforge.net/ for partial annotations, and extensive manual parsing carried out by one annotator using a custom-developed visual tree editor, Annotald [^12] https://github.com/Annotald/annotald.
2.3 Other Historical Corpora
Pettersson and Borin (2019) provides a comprehensive survey on existing diachronic and historical corpora. The work by Keersmaekers and Van Hal (2024) presents a case study demonstrating how large-scale automated parsing of Greek papyri can create richly annotated diachronic resources. Chen and Liu (2025) have created a Chinese corpus from the last 30 years of news articles on land usage. Even the corpora in higher-resourced languages such as DIAKORP Kučera et al. (2015) (Czech), ARCHER Biber et al. (1994), and  Davies (2012) [^13] https://www.english-corpora.org/coha/ (English), and DTA Geyken et al. (2011) and GerManC Scheible et al. (2011) (German) differ in size, balance, annotation depth, and access models. DIAKORP offers seven centuries of Czech texts, though it lacks linguistic annotation. ARCHER samples English registers across four centuries in 50-year intervals, while COHA spans two centuries of American English with lemmatisation and POS tagging.
Penn Parsed Corpora of Historical English Taylor and Kroch (1994) (PPCHE) and SRCMF Stein and Prévost (2013) (Old French) are similar to FarPaHC in the sense that they, too, are manually annotated corpora which provide syntactic analyses suitable for structural studies. PPCHE, in particular, has influenced corpora in other languages through its Penn-Helsinki annotation scheme, facilitating cross-linguistic comparison. Similarly, the PROIEL treebank family Eckhoff et al. (2018) extends such comparisons to some Indo-European languages via aligned New Testament translations.
ReM Klein and Dipper (2016) (Middle High German), RIDGES Odebrecht et al. (2017) (German-Science), and the Swedish Culturomics Gigaword corpus Eide et al. (2016), offer layered annotation or harmonised spellings for OCR quality control. While all corpora use some metadata scheme to provide critical contextual information such as date, genre, region, and authorship, beyond that, the metadata coverage varies widely. However, it can be noted that TEI-based [^14] Text Encoding Initiative (TEI) guidelines metadata schemes are popular among European language corpora.
3 Methodology
In this section, we describe the methodology used to create this dataset from the ground up. The process involved careful attention to detail at every stage, from planning to the final presentation, ensuring that the data is valid and of high quality. The procedure included addressing copyright laws in Sri Lanka, acquiring data, extracting text, and performing post-processing and formatting of the data as shown in Figure 1.
Figure 1: Summary of the Methodology Used in the Creation of SiDiaC.
3.1 Dataset Assembly
At first, we began acquiring Sinhala literature, including both fiction and non-fiction books, from the Internet Archives. However, the amount of data we were able to gather was quite limited. As a result, we decided to turn to the primary institution dedicated to Sinhala literature: the National Library (Natlib) of Sri Lanka [^15] https://www.natlib.lk/, which has its own digital repository [^16] https://diglib.natlib.lk/.
In the digital repository, we were able to organise all available content chronologically by issue date, allowing us to see publications printed dating back to 1800 CE. We carefully selected the book title, author name, identifier number, and collection name for each book from that point onward. This process required careful filtering, as most of the available content consisted of gazettes and police reports.
We identified 233 unique books printed between 1800 CE and 1955 CE in the Natlib digital repository (this is based on the issued date, not the written date). Of these, only 12 books were available for open access; we had to request access to the remainder. Our initial plan was to collect 100 sentences per year. To achieve this, we estimated that obtaining five pages of text from each book, excluding content pages and the preface, would provide us with more than 100 sentences, which amounts to approximately 1500 to 2000 word tokens, assuming there are about 15 to 20 words per sentence. Therefore, for the closed-access books, we requested five pages from each one. The process of obtaining access to these books was very difficult because most of them were part of the Rare Books Collection.
3.2 Annotation by Written Date
The issue date of the identified books was clearly stated in the digital repository at Natlib. However, this does not imply that the books were actually written during those specified dates. In fact, a book could have been written centuries earlier, while the printed version was released much later. Therefore, a comprehensive analysis was conducted to ensure that the written year of each book was accurately represented, ensuring that the resulting SiDiaC dataset accurately represents a proper diachronic corpus.
Upon the recommendation of experts in Sinhala linguistics, we identified a comprehensive book on Sinhala literature that claims to encompass literature information from its inception until 1994 CE Sannasgala (2015). This text served as the primary reference for the establishment of the respective date anchors, employing both time periods and specific years as outlined. The date ranges identified in Sannasgala (2015) correspond either to the period during which the book was authored or to the time period in which the author lived.
The process of determining the written dates for the books became more complicated because some books in the dataset include commentaries and discourses on earlier works. In this version of the dataset, these cases are tied to the original earlier book’s written date, as they contain both the information from the original book and its corresponding commentary (often written centuries prior, with extensive sections given as direct quotes without paraphrasing).
3.3 Challenges from Copyright Laws
During the planning stage, one of the biggest challenges we faced was managing copyright issues. To address this, we conducted a thorough analysis of copyright laws in Sri Lanka, which are governed by the Intellectual Property Act No. 36 of 2003171717https://www.gov.lk/wordpress/wp-content/uploads/2015/03/IntellectualPropertyActNo.36of2003Sectionsr.pdf.
According to this act, copyright in Sri Lanka is generally protected for the life of the author, plus an additional 70 years after their death. In cases where the author is unknown, copyright protection lasts for 70 years from the date of first publication. As a result, we focused on literature where the author passed away before 1955, as well as works by unknown authors that were published before 1955.
3.4 Data Filtration
We initially identified 233 unique books, but after careful consideration of several factors, we ultimately selected only 46. Our selection process was influenced by the availability of scanned copies, the written dates of the works, and compliance with copyright laws as illustrated in Figure 2.
Figure 2: Sequential Data Filtration Procedure
Our first limitation was the availability of scanned copies at the digital repository of Natlib, which restricted us to 221 books. Additionally, we were only able to determine the written dates or periods for 65 of the 233 books. Taking both the availability of scanned copies and the accessibility of written dates into account, the number of selected records was reduced to 59. We then further refined this selection following the copyright law based filtering process we discussed, resulting in a final total of 46 books as presented in Table 6.
3.5 Text Extraction using OCR
As the digital repository of Natlib shared the scanned copies of the requested books, we were forced to extract text information from the documents. Therefore, we selected an Optical Character Recognition (OCR) engine to get this task done.
In the comparative study conducted by Jayatilleke and de Silva (2025), five different OCR engines were analysed for text extraction using a synthetically created image-text dataset for Sinhala. Based on this study, we identified two standout OCR engines: Google Document AI [^18] https://cloud.google.com/document-ai/ and Surya [^19] https://github.com/VikParuchuri/surya, with Surya being reported to outperform all other systems compared. However, during our text extraction process, we found that, under realistic conditions (unlike the synthetic conditions used in the study), Google Document AI provided more accurate results.
Document AI is a service provided by Google Cloud Platform (GCP) [^20] https://cloud.google.com/. In this platform, we created a processor and utilised its Application Programming Interface (API) key to conduct OCR. The processor can handle a maximum of 15 pages at a time; however, this was not an issue since all of our scanned copies contained 5 to 8 pages, as shown in Figure 3. Throughout the procedure, we ensured that we obtained the model confidence for every page of each processed document. We then calculated the average confidence score, which is included in the metadata file of each book folder.
This OCR processor has demonstrated that it can perform text recognition that goes beyond simple extraction. It adapts effectively and generates words in modern Sinhala spelling while also taking into account morphology, where morphemes are formed accordingly, as explained in the section 4.1.
Figure 3: Distribution of Page Counts of Scanned Copies
3.6 Post-Processing Extracted Text
Although the OCR accuracy averaged 96.84% across all documents, the formatting issues were significant enough to require manual adjustments. The Document AI’s advanced performance helped to streamline manual post-processing tasks, significantly reducing the time required for that work.
The post-processing includes correcting the following text formatting issues:
  • Spacing Errors: This involves fixing incorrect or inconsistent spacing between words and sentences, such as missing spaces, uneven spaces and extra spaces.
  • Multi-column Text: This refers to texts containing errors where two columns were treated as a single column, resulting in entire horizontal lines being extracted without properly traversing each column separately.
  • Misplaced Words/Phrases: This addresses instances where words or phrases are out of order, leading to illogical text.
  • Paragraph and Line Indentation: This involves standardising the indentation of paragraphs and individual lines. This could mean adding consistent indents to new paragraphs (e.g., poem blocks) and removing incorrect indents.
  • Removal of Seal Context: This involves identifying and eliminating specific phrases or watermarks that represent seals or official stamps.
  • Page Number Removal: This focuses on identifying and deleting page numbers that appear within the body of the text, as they are part of the document’s structure rather than its content.
While language understanding was not a strict requirement for addressing these formatting-related factors, all manual post-processing procedures were carried out by the authors using a human-in-loop strategy Lamba and Madhusudhan (2023). This approach involved correcting formatting errors within a single window that contained both page scans and editable transcripts Christy et al. (2017). The authors responsible for these corrections are native Sinhala speakers. The post-processing steps applied, along with examples, are further discussed and illustrated in Appendix A.
An important finding of this study was the presence of Pali, Sanskrit, and minimal English in certain records of SiDiaC. The inclusion of Pali and Sanskrit can be attributed to the fact that most historical texts are related to religion. Notably, both Pali and Sanskrit are written in the Sinhala script in all cases. In this study, we chose not to remove the content in these languages to avoid losing context.
3.7 Creation of Metadata Files
The dataset consisted of folders, each dedicated to a specific book. Within each folder, there is a text file along with a metadata file. The metadata files contain information such as the title and author names in both Sinhala and romanised forms, as well as the genre, issue date, written date, and the OCR confidence level for each particular book. Most of these information fields were identified through the LatinlSE corpus McGillivray and Kilgarriff (2013). The overall composition of the SiDiaC corpus, including folder and file level examples, is illustrated in Figure 4.
Figure 4: Composition of the SiDiaC Corpus
The title of each book was consistently provided. When known, the authors’ names were included; if the authors were unknown, they were labelled as "unknown." The issued date corresponds to the published year as listed in the digital repository of Natlib, while the written date was determined by referring to Sannasgala (2015), as explained in the section 3.2.
The genres of the books were selected based on the details provided by Sannasgala (2015), as well as the content evaluated by authors who are native Sinhala speakers. The classification process occurs at two levels. The primary level is broad and divides the books into two categories: ‘Fiction’ and ‘Non-Fiction.’ The secondary level is more specific, categorising the content of the books into five distinct classes: religious, history, poetry, language, and medical. This approach was inspired by the methodologies followed in IcePaHC and DIAKORP Rögnvaldsson et al. (2012); Kučera et al. (2015) corpora to ensure diverse genres. It is important to note that the first level of categorisation was applied to all documents, while the second level of categorisation was applied only to the books that fell into the selected specific categories.
Furthermore, the average OCR confidence per page is included, as mentioned in the section 3.5. In addition, we romanised the titles and author names for each book, which involved transliterating Sinhala content into the Latin alphabet. An example of a metadata record is shown in Figure 5.
Figure 5: An example of a metadata record in SiDiaC
4 Analysis of SiDiaC
4.1 OCR Performance
As previously noted, the performance of Document AI extended beyond simple text extraction as shown in Figure 6. These improvements helped to streamline manual post-processing tasks, significantly reducing the time required for that work.
Two significant additional steps were clearly observable to be performed by Document AI.
  1. 1.
    Text Modernisation: The historical development of the Sinhala language has resulted in various eras of syntax being linguistically equivalent but different in grapheme representation Nandasara and Mikami (2016). It is clear that Document AI adjusts to generate text in modern Sinhala syntax, ensuring a consistent and unified syntax throughout the dataset. Given that the change is only at the grapheme level, this does not violate the syntactic or semantic properties of the word.
  2. 2.
    Morpheme Segmentation: In Historical Sinhala, certain words are combined without spaces, forming a closed compound Gaikwad and Saini (2024). This phenomenon was accurately identified by Document AI, which effectively performs morpheme segmentation.
Figure 6: Examples of Sinhala Text Modernisation and Morpheme Segmentation in Document AI
Some characters in SiDiaC literature do not exist in the Sinhala Unicode. Therefore, mapping old characters to modern ones was an essential task performed by Document AI. Morpheme segmentation is crucial for maintaining consistency among words. If this issue is not addressed, combined multi-word expressions may be treated as unique words during the word embedding process, which can lead to significant differences in results when analysing semantic meaning.
4.2 Evaluation of Metadata
The dataset spans from the 5th to the 20th century CE, making it the longest continuous diachronic Sinhala corpus created to date. It covers many significant time periods, from the Anuradhapura era (377 BCE – 1017 CE) to just after Sri Lanka gained independence from Britain in 1948. This extensive timeframe allows for a representation of various changes in the language over the centuries.
Primary Category Secondary Category Total
Non-Fiction Fiction Religious History Poetry Language Medical
5th 1 0 0 0 0 0 1 1
13th 7 1 5 0 1 2 0 8
14th 2 2 3 1 0 0 0 4
15th 1 4 2 0 3 0 0 5
18th 3 0 1 0 0 1 1 3
19th 6 3 2 2 3 2 0 9
20th 12 4 5 2 5 3 0 *16
Total 32 14 18 5 12 8 2 46
Table 1: Distribution of Books Across Centuries and Genres. *The total count for the secondary category in the 20th century amounts to 15, while the overall number of books is 16. This discrepancy arises because the book ‘Hithopadhesha Sannaya’, which offers advice, was not classified under any of the five secondary categories.
Assuming that the books with specified date ranges are attributed to the upper bound year, an analysis of the number of books per century was conducted, as illustrated in Table 1. The analysis reveals that the distribution of books in the corpus is heavily skewed toward the 20th century, with 28 out of 46 records originating after the 18th century. This trend may largely be attributed to the introduction of the printing press to Sri Lanka by the Dutch in 1737, which thereafter popularised book printing in the country Wickremasuriya (1978); Nandasara and Mikami (2016).
In the first level of genre classification between fiction and non-fiction, it is evident that there are more non-fiction books than fiction books in the corpus. At the second level of genre classification, religious texts and poetry dominate among the five categories. This predominance is largely due to the close relationship between Sinhala literary culture and Theravada Buddhism, which provided both subjects and a framework for preserving texts. Additionally, the influence of Sanskrit kavya traditions and courtly patronage, which valued literary artistry and prestige, also played a significant role Hallisey (2003).
The author of the book is known for 32 out of 46, while the remaining books are labelled as “Unknown." Only three authors have published more than one book in this dataset: two authors each have two books, and one author has four.
The OCR confidence levels are extremely high, with an average of 96.84% across all books and a minimum confidence score of 85.53%. Despite these encouraging figures, it is clear that Document AI encountered various types of errors, which we largely addressed during the post-processing phase as discussed in Appendix A. The accurate identification of characters and words likely contributes to these strong confidence scores; however, most errors appear to arise from the challenges presented by complex content formats.
4.3 Evaluation of the Corpus
SiDiaC consists of 58,027 word tokens that were filtered using regex, retaining only Sinhala and Latin characters, and subsequently tokenised by whitespace. The corpus contains 833 words in Latin script, which accounts for just 1.42% of the entire dataset. Also, the complete dataset comprises 22,837 unique word tokens in Sinhala script, which accounts for 39.36% unique word coverage of all words. This total word token count, while not in the range of millions, such as the COHA corpus for English, comfortably passes the 53,000 token count of FarPaHC for Faroese, which is in the same language resource category as Sinhala according to Ranathunga and de Silva (2022).
Figure 7: Total vs Unique Word Counts per Century
In the 5th century, 72.98% of words were unique, while the 13th century had 52.12% unique words. The 14th century saw 54.1%, the 15th century 55.97%, and the 18th century 55.05%. The 19th century featured 54.57%, but by the 20th century, the percentage dropped to 44.49%. As illustrated in Figure 7, generally a higher word count correlates with a lower percentage of unique words across the centuries.
Additionally, we conducted an analysis to identify the stopwords in the corpus by examining word tokens at the century level. Following the method described by Wijeratne and de Silva (2020) for their contemporary Sinhala corpus, we used a combination of word frequency analysis and manual vetting to identify appropriate stopwords from the corpora. To achieve this, we first counted the frequency of each unique word for each century and then converted these frequencies into z-scores.
Next, we calculated $\infty<Z<6.1027$ for a 99.80% threshold. In the 20th century, we observed the highest number of word tokens, which was used to establish the threshold, assuming that it provided adequate coverage of stop words throughout the entire corpus. After sorting the words in the 20th century by z-score in descending order, we manually inspected the words to determine the ideal z-score as the upper limit.
Over the centuries, the number of identified stopwords has varied accordingly. In the 5th century, there were 47 stopwords, followed by 42 in the 13th century, 39 in the 14th century, 28 in the 15th century, 44 in the 18th century, 65 in the 19th century, and finally, 61 in the 20th century.
The stopword that ranked highest across six of the seven centuries was ‘ [Uncaptioned image] ’(was, became), while ‘ [Uncaptioned image] ’(or, either) ranked highest in the fifth century. Additionally, other frequently occurring stopwords include [Uncaptioned image] (that, if), ‘ [Uncaptioned image] ’(that[named]), ‘ [Uncaptioned image] ’(that, those), ‘ [Uncaptioned image] ’(this, these), ‘ [Uncaptioned image] ’(or, interrogative particle), ‘ [Uncaptioned image] ’(while), ‘ [Uncaptioned image] ’(by), and ‘ [Uncaptioned image] ’(be [exists]). However, it was observed that certain words, such as ‘ [Uncaptioned image] ’(honorific for a revered person) and ‘ [Uncaptioned image] ’(Quality) were also receiving high z-scores, potentially due to the corpus’s strong connection to Buddhist literature. The methodology for identifying the top 48 stopwords in the entire corpus and their depiction is presented in Appendix C.
5 Conclusion
In this study, we introduced SiDiaC, a diachronic corpus of the Sinhala language. The corpus contains approximately 58k tokens, categorised into genres at two levels, and spans from the 5th century to the 20th century. This makes it the first diachronic Sinhala corpus ever created, which can serve as a foundational dataset for enhancing historical corpora in the Sinhala language. The entire process involved carefully identifying literature from the Natlib of Sri Lanka, followed by data filtering, date annotation, text extraction from PDF images, and post-processing. We also created metadata files containing important information about each book.
The complete corpus was thoroughly analysed, highlighting the powerful OCR performance of Document AI beyond simple text extraction. This was followed by a detailed evaluation of the dataset based on the metadata of all the books. Additionally, a comprehensive analysis was conducted at the word token level to ensure the identification of important findings within the corpus.
In future studies, this dataset can be further enhanced by using more literary works and improving its quality through better post-processing techniques. It provides capabilities to study language evolution, including lexical semantic change, the appearance of new words (neologisms), and corpus-based lexicography, among other applications.
Limitations
The creation of the corpus went through different types of limitations due to various challenges we faced.
Literature Identification: While we recognised the Department of National Archives [^21] https://websnew.lithium.lk/archives/ of Sri Lanka also as a credible source, data acquisition was conducted only from the National Library of Sri Lanka due to permission constraints.
Data Filtration: Out of the 221 scanned copies acquired, we were able to identify the written dates or periods for only 59 of them. The written dates of the books were annotated based on the lifespans of well-known authors, while the majority of the remainder were annotated relying heavily on the work by  Sannasgala (2015), which represents an overreliance on a single source.
Post-Processing after OCR: Under this process, corrections were initiated to address identified formatting issues, while possible identification errors at the word or character level were not addressed.
Code-Mixed Data: It was noted that the corpus contains code mixing of Pali, Sanskrit, and English languages, but the removal of text in these languages from the corpus has not been done.
Commentary Books: The identified books primarily named with the term ‘ [Uncaptioned image] ’ (meaning commentaries) will include two written dates for the original and the commentary. However, these instances were anchored to the well-known original version without removing the commentary.
Appendix A Post-Processing Extracted Text
During this phase, we addressed six types of formatting issues. This careful task was carried out by the authors of this study, who are native Sinhala speakers.
Certain literary works contained unwanted text referred to as seal context, which did not relate to the books’ actual content. As a result, this text was identified and removed from the book files in the dataset. Some examples of these seal contexts can be seen in Table 2.
Table 2: Examples of seal contexts found in the PDF image documents and the corresponding extractions by Document AI that were removed during post-processing.
Table 3: Examples of spacing errors after OCR using Document AI on images and their corresponding corrections. *Note that the OCR extractions depicted were not exact; some final words were completely unidentified, which were added manually, and some had line breaks in awkward places.
Table 4: Examples of errors in multi-column text after OCR using Document AI on images and their corresponding corrections. *Note that the OCR extractions shown here are not exact, as we could not fully represent an entire page that experienced this type of error in real case scenarios.
[…]
It was notable to see errors in spacing in poems, especially when the last word or letter of a word is separated by multiple spaces that were not correctly detected by Document AI, as shown in Table 3. Although correcting these errors may not hold significant semantic importance, accurately replicating the structure of the poems is crucial for preserving the original form of the books in case any downstream task using our dataset requires it. But for any task that does not require the original structure and would work on the lexical or semantic properties of the writing, the Document AI output is adequate.
There were also multi-column texts, particularly evident in poetry books such as ‘Kavya Wajrayudhaya - Palamu Kotasa’. These texts contained errors where two columns were treated as a single column, resulting in entire horizontal lines being extracted without properly traversing each column, as shown in Table 4. To address these errors, corrections were made by mimicking the original structure of the books to preserve their intended format. The corrections provided in the table demonstrate that OCR outputs can be restructured faithfully only through column-aware preprocessing (for example, layout analysis, region detection, or image segmentation).
The misplaced words and phrases appeared multiple times due to varying spacing and text positioning styles used in different books. Unlike simple character substitutions, these errors dislocate words or phrases from their expected syntactic and semantic slots. For example, text fragments are merged across lines, while key markers or paragraph boundaries vanish. This was particularly noticeable at the beginning of the first paragraph in texts with drop caps. Similarly, in poetry, irregular line breaks caused by extra spaces before the last letter or word were among these issues. A misplaced suffix or dislocated phrase could invert rhetorical emphasis or obscure reference chains. Importantly, these errors also impact computational parsing, where algorithms expecting consistent lineation and phrase boundaries will misinterpret discourse structure. These issues are illustrated in Table 5. Another one of the most common issues addressed during post-processing was the indentation errors in paragraphs and lines. These errors were often found in the first line of paragraphs, with certain poems being misaligned, and page headings being centred incorrectly. A few examples of these issues are presented in Table 5.
Document AI also captured meta information on the physical book such as page numbers. During the text extraction, these were removed from the books because they do not add any value to the presented corpus. These numbers were typically placed at the top or bottom, and they were centred or right-aligned in different literary works.
[…]
Primary Secondary
Adhimasa Dheepanaya Madhampe Dhammathilaka Himi Non-Fiction Religious 1896 1850 - 1896 0.9984
Adhimasa Winishchaya Walikande Sri Sumangala Himi Non-Fiction Religious 1904 1850 - 1904 0.9971
Adhimasa Sangrahawa Madhampe Dhammathilaka Himi Non-Fiction Religious 1903 1850 - 1903 0.9692
Anagathawanshaya: Methe Budu Siritha
Watadhdhara Medhanandha Himi;
Siri Parakumabahu;
Wilgammula Sangaraja Himi
Fiction Religious 1934 1325 - 1333 0.9992
Ashoka Shilalipi saha Prathimakarana Winishchaya D. E. Wickramasuriya Non-Fiction History 1919 1916 0.9989
Okandapala Sannaya hewath Balawathara Liyana Sanna Don Andhris Silva Non-Fiction Language 1888 1760 - 1778 0.9884
Kavya Wajrayudhaya - Palamu Kotasa Engalthina Kumari Fiction Poetry 1889 1825 - 1893 0.9254
Kavyashekaraya Thotagamuwe Rahula Himi Fiction Poetry 1872 1408 - 1491 0.9813
Kudusika Unknown Non-Fiction Poetry 1894 1270 - 1293 0.9988
Kusajathaka Wiwaranaya (Prathama Bagaya) Munidhasa Kumarathunga Non-Fiction Religious 1932 1887 - 1932 0.9701
Gadaladeni Sannayai Prasidhdha wu
Balawathare Purana Wyakyanaya
Hikkaduwe Sri Sumangala Himi Non-Fiction Language 1877 1827 - 1911 0.9970
Jubili Warnanawa John de Silva Non-Fiction Language 1887 1857 - 1922 0.9957
Dhaham Sarana Unknown Fiction Religious 1931 1220 - 1293 0.9891
Dhaladha Pujawaliya Unknown Non-Fiction History 1893 1325 - 1333 0.9978
Dhurwadhi Hardhaya Widharanaya Sri Dhanudhdharacharya Non-Fiction Religious 1899 1854 - 1899 0.9919
Dhampiya Atuwa Gatapadaya D.B. Jayathilaka Non-Fiction Religious 1932 1868 - 1932 0.9241
Dharma Pradheepikawa hewath
Mahabodhiwansha Parikathawa
Unknown Non-Fiction Religious 1906 1187 - 1225 0.9682
Dharmapradeepikawa Gurulu Gomeen Non-Fiction Religious 1951 1187 - 1225 0.9786
Nikam Hakiyawa Munidhasa Kumarathunga Fiction Poetry 1941 1887 - 1941 0.8932
Nikaya Sangrahaya hewath Shasanawatharaya
Unknown
Non-Fiction Religious 1922 1390 0.9754
Nidhahase Manthraya
S Mahinda Himi
Non-Fiction Poetry 1938 1901 - 1938 0.8997
Pansiya Panas Jathaka Potha
Unknown
Fiction Religious 1881 1303 - 1333 0.9987
Parawi Sandheshaya
Unknown
Fiction Poetry 1873 1430 - 1440 0.9902
Parani Gama
Galpatha Kemanandha Himi
Non-Fiction History 1944 1944 0.9846
Budhdha Sikka hewath Kudu Sika
Unknown
Non-Fiction History 1898 1270 - 1293 0.9699
Mage Malli
G. H Perera
Non-Fiction Poetry 1938 1886 - 1938 0.8659
Mahawansa Teeka
Hikkaduwe Sri Sumangala Himi
Non-Fiction History 1895 1827 - 1895 0.9978
Muwadew da Wiwaranaya
Munidhasa Kumarathunga
Non-Fiction Religious 1949 1887 - 1944 0.8710
Moggalalayanawyakaranan
Moggallana Himi
Non-Fiction Language 1890 1070 - 1232 0.9051
Moggallana Panchika Pradeepaya
Unknown
Non-Fiction Language 1896 1070 - 1232 0.9918
Liyanora Nadagama
Unknown
Fiction Poetry 1936 1852 - 1927 0.9935
Wibath Maldhama
Kirama Dhammarama Himi
Non-Fiction Language 1906 1821 0.9986
Waidya Chinthamani Baishadhya Sangrahawa
Unknown
Non-Fiction Medical 1909 1706 - 1739 0.9965
Wyakarana Wiwarana hewath
Sinhala Bashawe Wyakaranaya
Munidhasa Kumarathunga
Non-Fiction Language 1937 1887 - 1937 0.9029
Sadhdharma Rathnawaliya -Prathama Bagaya
Dharmasena Himi
Non-Fiction Religious 1930 1220 - 1293 0.9962
Sanna sahitha Abhisambodhi Alankaraya
Waliwita Saranankara Sangaraja Himi
Non-Fiction Religious 1897 1698 - 1778 0.9989
Sanna sahitha Salalihini Sandheshaya
Unknown
Fiction Religious 1859 1450 0.9909
Sanskrutha Shabdhamalawa
hewath Sanskrutha Nama Waranagilla
Rathmalane Dharmaloka Himi
Non-Fiction Language 1876 1828 - 1887 0.9671
Sarartha Sangrahawa: Prathama Bhagaya
Srimadh Budhdhadhasa Rajathuma
Non-Fiction Medical 1904 398 - 426 0.9997
Sithiyam sahitha Mahiyangana Warnanawa
Unknown
Fiction Poetry 1898 1878 0.9989
Sithiyam sahitha Sadhdharmalankaraya
Unknown
Non-Fiction Religious 1954 1398 - 1410 0.9810
Sithiyam sahitha Siyabas Maldhama
Kirama Dhammanandha Himi
Fiction Poetry 1894 1820 0.9256
Sithiyam sahitha Sinhala Mahawanshaya
D.H.S Abhayarathna
Non-Fiction History 1922 1874 0.9549
Sinhala Wyakaranaya enam Sidath Sangarawa
Hikkaduwe Sri Sumangala Himi
Non-Fiction Language 1884 1827 - 1911 0.9886
Hansa Sandheshaya
C.E. Godakumbure
Fiction Poetry 1953 1457 - 1465 0.8553
Hithopadhesha Sannaya
Waligama Sri Sumangala Himi
Non-Fiction
- 1884 1825 - 1905 0.9871
Table 6: The metadata information for all the literature used in the creation of this dataset.
Appendix B Utilized Literary Works
During this study, we collected 46 literary works from the Natlib of Sri Lanka, including 32 by recognised authors and the rest by ’unknown’ authors. Munidhasa Kumarathunga contributed four books, Madhampe Dhammathilaka Himi and Hikkaduwe Sri Sumangala Himi contributed two each, and the remaining authors each had one book, totalling 27 unique authors.
The metadata includes the title in Sinhala, the romanised title, the author’s name, the romanised author’s name, the genre, the issue date, the writing date, and the OCR confidence level. The complete metadata for each literary work used in this compilation of the dataset can be found in Table 6 with the titles and authors’ names presented in romanised Sinhala. This diachronic spread ensures coverage of Sinhala evolution across medieval, pre-modern, and modern stages. Religious texts dominate, reflecting both preservation biases and the centrality of Buddhism in Sinhala literary culture.
Stopword Meaning [in Context] 5th 13th 14th 15th 18th 19th 20th
[Uncaptioned image] [that which came to] be X X X X X X X
[Uncaptioned image] that, if X X X X X X X
[Uncaptioned image] that [named] X X X X X X X
[Uncaptioned image] that, those X X X X X
[Uncaptioned image] this, these X X X X X X X
[Uncaptioned image] or, interrogative particle X X X X X X
[Uncaptioned image] while X X X X X X
[Uncaptioned image] by X X X X X X
[Uncaptioned image] be [exists] X X X X X X
[Uncaptioned image] that [which was] X X X X X X
[Uncaptioned image] be [affirmed] X X X X
[Uncaptioned image] or X X X
[Uncaptioned image] and X X X X X X
[Uncaptioned image] by X X X
[Uncaptioned image] is [so named] X X X X
[Uncaptioned image] with, by X X X
[Uncaptioned image] be [affirmed] X X X
[Uncaptioned image] * Honourable [suffix] X X X
[Uncaptioned image] do [end of action] X X X
[Uncaptioned image] to X X
[Uncaptioned image] because X X X X
[Uncaptioned image] * Quality X
[Uncaptioned image] * Great [Prefix] X X X X
[Uncaptioned image] [the fact] that X X X
[Uncaptioned image] so [called] X X X X
[Uncaptioned image] be [reported] X X X
[Uncaptioned image] is [reported] X X X
[Uncaptioned image] by [means of] X X X
[Uncaptioned image] while, when X X X
[Uncaptioned image] like [in manner of] X X
[Uncaptioned image] [poetic suffix] X X
[Uncaptioned image] be [affirmed] X X
[Uncaptioned image] be [reported] X X
[Uncaptioned image] * Great [Prefix] X X
[Uncaptioned image] [poetic suffix] X X
[Uncaptioned image] be [known to exist] X X
[Uncaptioned image] at, after [so declared] X X
[Uncaptioned image] a/the/one X X
[Uncaptioned image] him X X
[Uncaptioned image] not X X
[Uncaptioned image] also [in similar manner] X X
[Uncaptioned image] therein X X
[Uncaptioned image] for X
[Uncaptioned image] * Buddha X X
[Uncaptioned image] * Quality X X
[Uncaptioned image] [after] be X X
[Uncaptioned image] herein X
[Uncaptioned image] do [end of action] X X
Table 7: The top 48 stopwords and their presence in each century after applying the threshold $\infty<Z<6.1027$. *Note that some words, which are not technically stopwords, were included in this analysis. This may have occurred due to the limited availability of literary works in certain centuries and a bias toward Theravada Buddhism in the selected literature.
Appendix C Analysis of Stop Words in SiDiaC
The stopword analysis was conducted using the z-score calculation, as detailed in section 4.3. During this analysis, we identified a union set of words that exceeded the established threshold, resulting in a total of 194 unique words. The words were sorted by their average z-score for each century in descending order. The list of words that were above the set threshold was cross-checked with the union set, illustrating their availability in each century. The top 48 words with the highest mean z-scores are displayed in Table 7.
The continued inclusion of particles and suffixes shows continuity of core grammatical function words. Their presence across the 5th–20th centuries suggests remarkable diachronic stability in Sinhala morphosyntactic scaffolding. However, the table also reveals anomalies: certain lexical items not traditionally classified as stopwords appear in the stopword list. This distortion is likely due to corpus composition (religious texts with Buddhist themes dominate some centuries, inflating the relative frequency of doctrinal terms). Another observation is the persistence of poetic suffixes in older centuries, gradually tapering in modern texts. This diachronic shift may point to a movement away from poetry and towards prose. The presence/absence patterns also reveal data sparsity in some centuries, as marked by gaps where certain stopwords do not appear due to limited surviving texts. To wit, it is as interesting (mayhap more) to note what is missing in the 20th century, given that it seems to count a majority of the candidate words among its stopwords. Note how the archaic forms of honorifics and some direct references to Buddhism have dropped out of the list. In earlier centuries, the preponderance of monastic authorship and the dominance of canonical or exegetical works meant that words indexing reverence and religious entities were unavoidable high-frequency items. Their disappearance, or at least their reduced prominence, in the most recent century may be reflecting how the subjects covered in the text have shifted from esoteric religious communication to comparatively more secular discourse.